# CC-Diff: Enhancing Contextual Coherence in Remote Sensing Image Synthesis

### v2

@@ 4.3 Foreground & Background Generation
Figure 3: Qualitative L2I results on DIOR-RSVG (top two rows) and DOTA (bottom two rows). CC-Diff not only synthesizes realistic foregrounds with accurate positioning but also generates more detailed backgrounds with stronger coherence to the foreground. Please zoom in for better details.
Foreground Instace Synthesis. As illustrated in Figure 2, the Conditional Generation Module synthesizes each foreground instance via distinctive branches in parallel. The input of each branch, termed foreground embeddings $\bm{h^{fg}_{i}}$, can be expressed as
$\bm{h^{fg}_{i}}=\left[\bm{h_{i}},\operatorname{CLIP}\left(c_{i}\right),\operatorname{MLP}\left(\operatorname{Fourier}\left({b_{i}}\right)\right)\right]$ (5)
which are exactly foreground tokens $\bm{h_{i}}$ concatenated with the embedding of class label $c_{i}$ and bounding box $b_{i}$. These input variables control the synthesized foreground instances also via cross-attention, which can be formulated as
$\bm{R_{i}^{fg}}=\operatorname{Softmax}\left(\frac{{Q{K_{i}^{fg}}}^{T}}{\sqrt{d}}\right){V_{i}^{fg}}\cdot{M_{i}}$ (6)
where ${K_{i}^{fg}}$ and ${V_{i}^{fg}}$ are projections of the foreground embedding $\bm{h^{fg}_{i}}$ and $Q$ is computed from the image feature map of the diffusion model backbone. The output of each cross-attention block is then regulated via an instance mask $M_{i}$. Notably, unlike most existing methods adopting binary masks, we resort to a non-uniform implementation using the Sigmoid() function, which can be written as
$\operatorname{Sigmoid\left(x,y\right)=\frac{1}{1+\operatorname{exp}\left(-1+\frac{{\left(x-{\mu}_{1}\right)}^{2}}{{\sigma}_{1}^{2}}+\frac{{\left(y-{\mu}_{2}\right)}^{2}}{{\sigma}_{2}^{2}}\right)}}$ (7)
where $\mu_{1}$, $\mu_{2}$ represent the center of the oriented bounding box and $\sigma_{1}=w/2$, $\sigma_{2}=h/2$. This implementation achieves smooth translation across instances and surrounding texture, as well as a convenient and intuitive representation of oriented objects typically in RS images.
Conditional Background Generation. For synthesizing background texture, besides the background queries $\bm{q^{bg}}$ encoding the low-level information of reference images, the relating queries $\bm{q^{rel}}$ are also involved to introduce contextual knowledge. Moreover, unlike existing approaches where such conditions also interact with the image feature through the cross-attention layer, we propose to straightforwardly reinforce the awareness of foreground features by introducing the fusion of grounded feature maps of all instances, i.e., $\bm{R^{fused}}=\sum_{i=1}^{N}\bm{R^{fg}_{i}}$, into the cross-attention mechanism, which can be written as
$\bm{{R^{bg}}}=\operatorname{Softmax}\left(\frac{{Q{K^{bg}}}^{T}}{\sqrt{d}}\right){V^{bg}}$ (8)
Similar to the synthesis of foreground instances, $K^{bg}$ and $V^{bg}$ are projections obtained from the background embedding $\bm{h^{bg}}=\left[\bm{q^{bg}},\operatorname{CLIP}\left(\mathcal{P}\right)\right]$, which combines the learned background query $\bm{q^{bg}}$ with the embedding of global semantic $\mathcal{P}$. However, $Q$ is derived from $\bm{R^{fused}}$, rather than the image feature maps used for foreground synthesis as in Eq 6. In the end, feature maps of foreground and background are added up and sent to the decoder network for producing the final output images.
@@
CC-Diff (Ours) $\bm{21.78}$ $32.43$ $\bm{68.03}$ $\bm{67.68}$ $\bm{41.37}$ CC-Diff (Ours) $\bm{22.42}$ $30.90$ $\bm{47.42}$ $\bm{44.54}$ $\bm{22.21}$
5 Experiments
@@ 5.1 Experimental Settings
Datasets. Following datasets are used in experiments:
@@
CC-Diff (Ours) 52.62 77.51 58.09 37.57 63.18 38.05
Implementation Details. The pre-trained Stable Diffusion V1.4 [32] is used as the backbone for fine-tuning in all experiments. The global text prompt for describing the layout of each image is constructed following the rule-based protocol in [20]. All images are resized to $512\times 512$  [^1] Images in DOTA are evenly segmented into the same resolution, without considering the integrity of whole objects., with a maximum of 6 objects per image, consistent with recent L2I studies [59, 58]. For all experiments, CC-Diff is trained for 100 epochs with a batch size of 320 on 8 $\times$ NVIDIA A800 GPUs, with a fixed learning rate of $1e^{-4}$.
Benchmark Methods. We use generation methods from both RS and natural images as benchmarks. Txt2Img-MHN [49] and DiffusionSAT [12] are recent T2I methods for RS images, which are prompted by the previously mentioned rule-based text description. To ensure a fair comparison of spatial controllability and evaluate CC-Diff’s generalizability, we also include state-of-the-art L2I methods for natural image generation: GLIGEN[16], LayoutDiffusion [57], and MIGC [59].
Evaluation Metrics. Following [16, 57, 59, 6], we evaluate the performance of CC-Diff across three main aspects:
  • Faithfulness: Generated images are expected to align with the provided prompt, with global and local semantic consistency assessed by the Global and Local CLIPScores [1]. The YOLOScore [17] computed by an off-the-shelf Faster R-CNN detector [31] further evaluates the alignment of layout for input and output.
@@ 5.2.1 Qualitative Results on RS Datasets
Figure 3 presents a comparison of generated RS images. CC-Diff not only synthesizes realistic foreground instances at varying scales, with accurate positioning and orientation consistent with the given layout, but also generates backgrounds with more intricate textures. More importantly, it establishes a significantly more coherent and plausible relationship between the foreground and background. For example, in the first two cases, CC-Diff successfully renders the road going through the foreground instances, showing a reasonable association with the presence of the expressway service areas and toll station.
As for the benchmark methods, Txt2Img-MHN and DiffusionSat struggle to control the location of foreground objects, despite the inclusion of spatial information in global text prompts. Among the L2I approaches, MIGC achieves the best visual quality and semantic consistency. However, compared to CC-Diff, its rendered background lacks coherence with the foreground, resulting in an implausible landscape. Please refer to Supp. Mat. Sec 10 for more results.
@@ 5.2.2 Quantitative Results on RS Datasets
Visual Fidelity. CC-Diff achieves leading FID scores of 68.03 on DIOR-RSVG and 47.42 on DOTA, outperforming the second-best methods by 10.13 points (DiffusionSat: 78.16) and 16.53 points (MIGC: 63.95), respectively. This clear performance edge can be attributed to CC-Diff’s effective synthesis of realistic instance textures and its enhanced alignment with the overall scene context.
Semantic Faithfullness. CC-Diff demonstrates state-of-the-art performance in global and regional semantic alignment, as indicated by CLIPScore values. While DiffusionSat achieves a marginally higher Global CLIPScore (32.68 vs. 32.43), this is likely due to its larger RS-specific pre-training dataset [12]. Additionally, CC-Diff shows a clear advantage in YOLOScore, confirming that instances are well recognized by the object detector, with accurate layout retention reflected in the high mAP values.
Trainability. Following the data enhancement protocol in [6], we double the training samples using layout-based synthesis and assess detection accuracy with the expanded dataset. As shown in Table 2, CC-Diff consistently delivers the highest accuracy gains, improving mAP by 2.45 on DIOR-RSVG and 2.04 on DOTA. Please refer to Supp. Mat. Sec 11 for additional results in fine-grained settings. To the best of our knowledge, this is the first study to confirm the effectiveness of synthetic samples in improving object detection performance in RS.
5.3 Natural Image Synthesis
@@ 5.3.1 Qualitative Results on COCO
Table 4: Trainability ($\uparrow$) comparison on COCO. ’Baseline’ denotes accuracy with the unaugmented dataset.
@@ 5.3.2 Quantitative Results on COCO
We follow the evaluation protocol on RS datasets and assess the performance of CC-Diff on COCO by comparing it to benchmark methods.
Visual Fidelity. As shown in Table 3, CC-Diff achieves an FID score of 30.35 on COCO dataset. Although it is slightly lower than the seminal study GLIGEN [16] by 1.66 FID, CC-Diff outperforms MIGC (34.31) by 3.96 and LayoutDiffusion (37.26) by 6.91. This demonstrates the promising generalizability of CC-Diff to natural image datasets with more diverse and complex foreground attributes.
@@
Table 5: Ablation on Dual Re-sampler and FG-aware Attention.
  Dual Re-sampler FG-aware Attention CLIPScore $\uparrow$ FID $\downarrow$ YOLOScore $\uparrow$
@@
$21.69$ $32.39$ $72.39$ $63.85$ $37.68$
$\bm{21.91}$ $32.38$ $71.17$ $65.86$ $39.25$
$21.74$ $\bm{32.43}$ $69.56$ $64.62$ $39.22$
$21.78$ $\bm{32.43}$ $\bm{68.03}$ $\bm{67.68}$ $\bm{41.37}$
@@ 5.4 Ablation Study
Sequential Generation Pipeline. The core innovation of CC-Diff is its sequential generation pipeline, which explicitly models the contextual alignment and semantic interdependence between foreground and background. The Dual Re-sampler and FG-aware Attention mechanisms are central to this pipeline, and their individual contributions are assessed by omitting each one in turn.
@@
LLM-assisted T2I. While synthetic RS images show promising trainability, realistic layout guidance still requires extensive annotation. To address this, we explore using Large Language Models (LLMs) to generate plausible foreground layouts from the same rule-based textual descriptions [20] that prompted Txt2Img-MHN and DiffusionSat in previous experiments (Sec. 5.2).
As shown in Table 6, CC-Diff improves overall detection accuracy by 1.74 mAP and consistently achieves the highest performance gains across all trainability settings. Additionally, CC-Diff demonstrates strong global semantic consistency, reflected by a Global CLIPScore comparable to Ground Truth and an FID score that surpasses the second-best method by 6.87 points (77.48 to 70.61). Furthermore, Figure 5 demonstrates CC-Diff’s adaptability to LLM-generated layout conditions, showcasing its promising zero-shot capability and the diversity of this LLM-based T2I pipeline.
@@ 6 Conclusion
Supplementary Material
Due to space constraints in the main submission, we provide the detailed implementation and explanation of the proposed CC-Diff in this Supplementary Material. Specifically, we present the implementation details of CC-Diff, including the definition of the bounding box orientation, the rule-based protocol for constructing the global text prompt $\mathcal{P}$, the procedure for retrieving reference images for foreground and background synthesis from the dataset, and the specification of CC-Diff’s internal structure. Additionally, we include further experimental results to underscore the effectiveness of CC-Diff.
@@ 8 Construction of the Global Text Description and the GPT Prompt
9 Reference Image Retrieval
To improve the visual quality of synthesized results, we employ a method inspired by augmented image retrieval [], selecting reference images from the training set based on specific criteria to supply dense guiding information. The detailed procedure is outlined as follows:
  1. 1.
    We use SAM [13] to segment instances from patches cropped from training images based on HBB annotations. For each category, the segmented instances are sorted by resolution (in descending order), and the top-ranked ones (e.g., 200) are selected to build a reference dictionary. During inference, a reference instance, whose embedding is denoted as $\bm{f_{ins}}$ in the main submission, is randomly chosen from the dictionary based on the querying class label.
  2. 2.
    For the background reference image with embedding $\bm{f_{img}}$, we select the one that is semantically closest to the construction image caption $\mathcal{P}$ (described in Sec 8). To improve efficiency, we pre-compute and store the CLIP embeddings of the training images [27]. During retrieval, we compute the similarity between the CLIP embedding of $\mathcal{P}$ and the stored embeddings, then randomly sample one of the top five images with the highest similarity values as $\bm{f_{img}}$.
Figure 8 provides an illustration of the retrieved instances and reference images for further reference.
Ablation on Reference Image Guidance. We also assess the impact of integrating reference images retrieved from the training set. As shown in Table 7, using reference images for foreground synthesis notably improves the FID by 3.6 points (from 75.46 to 71.86) and slightly enhances the Local CLIPScore. Meanwhile, incorporating background reference images primarily contributes to higher YOLOScores and further improvements in FID.
@@ 11 Trainability AP
The detailed trainability results are provided in Tables 89, and 10, where the accuracy improvements for each individual object class are presented to facilitate a more thorough analysis.
From Table 8, it is evident that CC-Diff outperforms across multiple categories, particularly in complex and diverse scenarios. It achieves the highest scores in categories such as golf field (54.19), stadium (75.10), train station (28.87), harbor (12.26), storage tank (80.57), airport (47.57), and dam (36.40). These results underscore CC-Diff’s strong generalization capabilities and its effectiveness in handling a variety of scenarios within the DIOR-RSVG dataset, demonstrating its potential for remote sensing image analysis. This ability is further reflected in the DIOR dataset with layout prompts generated by GPT-4o (Table 9) and the DOTA dataset for RS object detection (Table 10).
Table 7: Ablation on reference, $\dagger$ indicates random sampling.
  FG Ref. BG Ref. CLIPScore $\uparrow$ FID $\downarrow$ YOLOScore $\uparrow$
Local Global $\text{mAP}_{50}$ $\text{mAP}_{50-95}$
$21.76$ $32.42$ $75.46$ $66.20$ $40.30$
$21.70$ $\bm{32.58}$ $71.90$ $67.10$ $40.40$
$21.80$ $32.32$ $71.86$ $65.30$ $38.60$
$\bm{21.81}$ $32.43$ $70.00$ $67.35$ $40.64$
$21.78$ $32.43$ $\bm{68.03}$ $\bm{67.68}$ $\bm{41.37}$
Figure 8: Detailed Dual Re-sampler Structure
Figure 9: Detailed Conditional Generation Module Structure
12 Detailed Implementation of CC-Diff
Dual Re-sampler As illustrated in Figure 8, Both BG-Resampler and FG-Resampler consist of four consecutive combinations of Cross-Attention (CA) and Feed Forward Network (FFN) layers. The CA layers in FG-Resampler enable the interaction between foreground queries ($\bm{q^{fg}}$) and the embedding of foreground instances ($\bm{f_{ins}}$ retrieved from the training set and then extract the foreground tokens ($\bm{h}$). To establish the connection between $\bm{q^{fg}}$ and $\bm{q^{bg}}$, relating queries ($\bm{q^{rel}}$) first interact with the grounded foreground tokens obtained through adding the bounding box formation into the foreground tokens through a GSA network and then concatenate with $\bm{q^{bg}}$. The combination results, together with the embedding of reference image ($\bm{f_{img}}$) are finally fed into the CA layers in BG-Resampler to obtain background tokens with overall coherence of the generated image.
Conditional Generation Module As illustrated in Figure 9, the Conditional Generation Module synthesizes each foreground instance via distinctive branches in parallel. The input of each branch, termed foreground embeddings $\bm{h^{fg}_{i}}$ are foreground tokens $\bm{h_{i}}$ concatenated with the embedding of class label $c_{i}$ and bounding box $b_{i}$. Specifically, a grounding MLP consisting of a Fourier Transform and a MLP network is utilized to encode $b_{i}$ into a positional embedding, and we use the CLIP text encoder to compute the embedding of the text description. These input variables control the synthesized foreground instances also via cross-attention, where the foreground embeddings $\bm{h^{fg}}$ interact with the image feature map of the diffusion model backbone. The output of each cross-attention block is then regulated via an instance mask $M_{i}$ which is implemented using the Sigmoid Function and then obtain the foreground features $\bm{R^{fg}_{i}}$.
Similar to foreground instances, the generation of background is controlled by the background embeddings ($\bm{q^{bg}}$) which combine the background tokens with the embedding of global semantic $\mathcal{P}$. Moreover, we straightforwardly reinforce the awareness of foreground features by introducing the fusion of grounded feature maps of all instances, i.e., $\bm{R^{fused}}=\sum_{i=1}^{N}\bm{R^{fg}_{i}}$, into the cross-attention mechanism to obtain background feature maps. In the end, feature maps of foreground and background are added up and sent to the decoder network for producing the final output images.

### v3

@@ 4.2 Dual Re-sampler for FG-BG Association
Figure 6: Qualitative L2I results on DIOR-RSVG (top three rows) and DOTA (bottom three rows). CC-Diff not only synthesizes realistic foregrounds with accurate positioning but also generates more detailed backgrounds with stronger coherence to the foreground. Please zoom in for better details.
4.3 CGM for FG & BG Generation
Given the condition embeddings $\{\mathbf{e^{fg}_{i}}\}_{i=1}^{N}$ and ($\mathbf{e^{bg}}$) from the Dual Re-sampler, the Conditional Generation Module (CGM) synthesizes grounded FG instances in parallel and renders the BG using their integrated representation. An illustration of CGM is provided in Figure 5.
FG Instance Synthesis. For the $i$-th FG instance, the condition embedding $\mathbf{e^{fg}_{i}}$ acts as the controlling signal for the synthesis process, and the rendered feature map $\mathbf{R_{i}^{fg}}$ is obtained via the CA mechanism as follows
$\mathbf{R_{i}^{fg}}=\text{CA}(\mathbf{f},\mathbf{e^{fg}_{i}})\odot{\mathbf{M_{i}}},$ (7)
where $\mathbf{f}$ denotes the incoming latent image feature to CGM from the previous layer in the LDM’s UNet, and $\mathbf{M_{i}}$ is the instance mask regulating the region of attention. Notably, rather than using a binary mask, we adopt a non-uniform mask derived from a rotated Sigmoid() function [44], enabling smooth transitions between instances and their surrounding textures (see Appendix Sec.9 for details).
FG-aware BG Rendering. Unlike most existing methods [41, 60, 62] that derive BG based on the image feature $\mathbf{f}$, we strengthen FG awareness by instead conducting BG synthesis based on the aggregation of all FG instance feature maps $\mathbf{R^{fused}}=\sum_{i=1}^{N}\mathbf{R^{fg}_{i}}$. This design promotes a tighter coupling of FG and BG at the feature level, resulting in smoother transitions in the generated image (see Sec. 5.4 for discussion). Specifically, we also compute the BG feature $\mathbf{R^{bg}}$ via the FG-aware Attention mechanism using CA, i.e., $\mathbf{R^{bg}}=\text{CA}(\mathbf{R^{fused}},\mathbf{e^{bg}})$, and finally add $\mathbf{R^{fused}}$ to $\mathbf{R^{bg}}$ to obtain the final output of CGM.
5 Experiments
@@
CC-Diff (Ours) $\bm{21.82}$ $32.36$ $\bm{70.68}$ $\bm{68.40}$ $\bm{41.92}$ CC-Diff (Ours) $\bm{22.60}$ $30.92$ $\bm{47.72}$ $\bm{45.09}$ $\bm{21.83}$
@@ 5.1 Experimental Settings
Datasets. Our experiments are conducted on two remote sensing (RS) datasets: DIOR-RSVG [57] with 17,402 images and the more challenging DOTA [49] with 2,806 images. We also use COCO2017 [20] to assess the generalizability of CC-Diff in handling diverse object categories and complex attributes in natural scenes. Please refer to Sec. 10 of the Appendix for details on dataset preparation.
@@
CC-Diff (Ours) 52.18 77.39 57.59 37.36 63.18 38.55
Implementation Details. We adopt the pre-trained Stable Diffusion V1.4 [33] as the backbone for fine-tuning in all experiments. The global text prompt for each image’s layout follows the rule-based protocol in [21], while all learnable queries are randomly initialized using a standard normal distribution. Following recent L2I efforts [62, 61], we resize all images to $512\times 512$  [^1] Images in DOTA are evenly segmented into the same resolution, without considering the integrity of whole objects. and allow a maximum of six objects per image. We train CC-Diff for $100$ epochs with a batch size of $320$ on 8×8$\times$8 × NVIDIA A800 GPUs, with a fixed learning rate of $1e^{-4}$.
Benchmark Methods. We select both remote sensing (RS) and natural image generation methods as benchmarks. For RS images, we adopt three recent controllable generation approaches: Txt2Img-MHN [52], DiffusionSAT [14], and AeroGen [41]. The first two rely on our previously described rule-based text prompts, while AeroGen additionally incorporates spatial layouts. For a fair comparison of spatial controllability and to assess CC-Diff’s generalizability, we also include three state-of-the-art layout-to-image (L2I) methods for natural image generation: GLIGEN [17], LayoutDiffusion [60], and MIGC [62].
Evaluation Metrics. Following [17, 60, 62, 7], we evaluate CC-Diff on three key criteria: (1) Fidelity, measured by the FID score [11] for perceptual quality; (2) Faithfulness, assessed using Global and Local CLIPScores [2] for semantic consistency and YOLOScore [18] for layout alignment; and (3) Trainability, evaluated by mean Average Precision (mAP) to test how well synthetic samples boost object detection accuracy. See Appendix Sec. 11 for further details.
@@ 5.2.1 Qualitative Results on RS Datasets
Figure 6 presents a comparison of generated RS images. CC-Diff not only synthesizes realistic foreground instances at varying scales, with accurate positioning and orientation consistent with the given layout, but also generates backgrounds with more intricate textures. More importantly, it establishes a significantly more coherent and plausible relationship between the foreground and background. For example, in the first three cases, CC-Diff successfully renders the road going through the foreground instances, showing a reasonable association with the presence of the expressway service areas, toll station and overpass.
As for the benchmark methods, Txt2Img-MHN and DiffusionSat struggle to control the location of foreground objects, despite the inclusion of spatial information in global text prompts. Among the L2I approaches, MIGC achieves the best visual quality and semantic consistency. However, compared to CC-Diff, its rendered background lacks coherence with the foreground, resulting in an implausible landscape. Please refer to Appendix Sec. 12 for more results.
@@ 5.2.2 Quantitative Results on RS Datasets
Visual Fidelity. CC-Diff achieves leading FID scores of $70.68$ on DIOR-RSVG and $47.72$ on DOTA, outperforming the second-best methods by $2.34$ points (GLIGEN: $73.02$) and $7.30$ points (AeroGen: $55.02$), respectively. This clear performance edge can be attributed to CC-Diff’s effective synthesis of realistic instance textures and its enhanced alignment with the overall scene context.
Semantic Faithfullness. CC-Diff demonstrates state-of-the-art performance in global and regional semantic alignment, as indicated by CLIPScore values. While DiffusionSat achieves a marginally higher Global CLIPScore ($32.68$ vs. $32.36$), this is likely due to its larger RS-specific pre-training dataset [14]. Additionally, CC-Diff shows a clear advantage in YOLOScore, confirming that instances are well recognized by the object detector, with accurate layout retention reflected in the high mAP values.
Trainability. Following the data enhancement protocol in [7], we double the training samples using layout-based synthesis and assess detection accuracy with the expanded dataset. As shown in Table 2, CC-Diff consistently delivers the highest accuracy gains, improving mAP by $2.01$ on DIOR-RSVG and $1.83$ on DOTA. Please refer to Appendix Sec. 13 for additional results in fine-grained settings.
5.3 Generalization to Natural Image Synthesis
@@ 5.3.1 Qualitative Results on COCO
Table 4: Trainability ($\uparrow$) comparison on COCO. ‘Baseline’ denotes accuracy with the unaugmented dataset.
@@ 5.3.2 Quantitative Results on COCO
Visual Fidelity. As shown in Table 3, CC-Diff achieves an FID score of $30.35$ on COCO dataset. Although it is slightly lower than the seminal study GLIGEN [17] by $1.66$ FID, CC-Diff outperforms MIGC ($34.31$) by $3.96$ and LayoutDiffusion ($37.26$) by $6.91$. This demonstrates the promising generalizability of CC-Diff to natural image datasets with more diverse and complex foreground attributes.
@@
Table 5: Ablation on Context Bridge and FG-aware Attention.
  Context Bridge FG-aware Attention CLIPScore $\uparrow$ FID $\downarrow$ YOLOScore $\uparrow$
@@
$21.68$ $32.33$ $72.30$ $63.08$ $37.49$
$\bm{21.90}$ $32.32$ $71.83$ $64.98$ $38.42$
$21.76$ $32.34$ $71.28$ $64.38$ $39.22$
$21.82$ $\bm{32.36}$ $\bm{70.68}$ $\bm{68.40}$ $\bm{41.92}$
Figure 8: Qualitative results from various CC-Diff variants. Incoherent regions marked by dashed yellow boxes and labeled on top.
@@ 5.4 Ablation Study
LLM-assisted T2I. While synthetic RS images demonstrate strong trainability, achieving realistic layout guidance still requires extensive annotation. To address this challenge, we leverage Large Language Models (LLMs) to generate plausible foreground layouts from the same rule-based textual descriptions [21] used to prompt Txt2Img-MHN and DiffusionSat in earlier experiments (Sec. 5.2).
As shown in Table 6, CC-Diff improves overall detection accuracy by $1.74$ mAP and consistently achieves the highest performance gains across all trainability settings. Additionally, CC-Diff demonstrates strong global semantic consistency, evidenced by a Global CLIPScore on par with Ground Truth and an FID score exceeding the second-best method by $6.10$ points ($76.71$ vs. $70.61$). Figure 9 further highlights CC-Diff’s adaptability to LLM-generated layout conditions, showcasing its promising zero-shot capability and the diversity enabled by this LLM-based T2I pipeline.
@@ 6 Conclusion
Appendix
Due to space constraints in the main submission, we provide the detailed implementation and explanation of the proposed CC-Diff in this Appendix. Specifically, we present the implementation details of CC-Diff, including the definition of the bounding box orientation, the rule-based protocol for constructing the global text prompt $\mathcal{P}$, the implementation of instance mask and the setting details of datasets and metrics used in the experiments. Additionally, we include further experimental results to underscore the effectiveness of CC-Diff.
@@ 8 Construction of the Global Text Description and the GPT Prompt
9 Implementation of the Instance Mask
In the process of FG instance synthesis, the output of each cross-attention block is regulated via multiplying by an instance mask $\mathbf{M}$ by element which is converted from bbox $\mathbf{b}=[x_{c},y_{c},w,h,\theta]$. Unlike most existing methods adopting binary masks, we resort to a non-uniform implementation using the Sigmoid() function, which can be written as
$\begin{split}\text{Sigmoid}(x,y)&=\frac{1}{1+\text{exp}(-1+\frac{{(x^{\prime}-{\mu}_{1})}^{2}}{{\sigma}_{1}^{2}}+\frac{{(y^{\prime}-{\mu}_{2})}^{2}}{{\sigma}_{2}^{2}})}\\ \begin{pmatrix}x^{\prime}\\ y^{\prime}\end{pmatrix}&=\begin{pmatrix}\cos{\theta}&\sin{\theta}\\ -\sin{\theta}&\cos{\theta}\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\end{split}$ (8)
where $\mu_{1}$, $\mu_{2}$ represent the center of the oriented bounding box ($x_{c},y_{c}$), $\sigma_{1}=w/2$, $\sigma_{2}=h/2$ and $x^{\prime},y^{\prime}$ are the results of rotating the grid coordinates ($x,y$) counterclockwise by $\theta$ angle.
10 Dataset Details
Following datasets are used in experiments:
11 Evaluation Metrics Setting
Following [17, 60, 62, 7], we evaluate the performance of CC-Diff across three main aspects:
  • Faithfulness: Generated images are expected to align with the provided prompt, with global and local semantic consistency assessed by the Global and Local CLIPScores [2]. The YOLOScore [18] computed by a YOLOv8 detector [13] further evaluates the alignment of layout for input and output.
@@ 13 Trainability AP
The detailed trainability results are provided in Tables 78, and 9, where the accuracy improvements for each individual object class are presented to facilitate a more thorough analysis.
From Table 7, it is evident that CC-Diff outperforms across multiple categories, particularly in complex and diverse scenarios. It achieves the highest scores in categories such as golf field ($54.65$), ground track field ($69.75$), train station ($28.63$), basketball court ($56.78$). These results underscore CC-Diff’s strong generalization capabilities and its effectiveness in handling a variety of scenarios within the DIOR-RSVG dataset, demonstrating its potential for remote sensing image analysis. This ability is further reflected in the DIOR dataset with layout prompts generated by GPT-4o (Table 8) and the DOTA dataset for RS object detection (Table 9).